OpenStudy (anonymous):

so i need help. its a circle equation. it says "find the value of x illistrated below" and the equation gives me 6x+10 with a 160 degree sign outside of the circle

OpenStudy (dumbcow):

not sure what you mean: the equation for a circle is (x-h)^2 +(y-k)^2 = r^2

OpenStudy (anonymous):

thats the illistration of it

OpenStudy (dumbcow):

oh thank you, geometry not circle equation

OpenStudy (anonymous):

oh lol k

OpenStudy (dumbcow):

remember the outside arc is always double the inside angle when vertex of angle is on outside of the circle half of 160 is 80 --> 6x+10 = 80 solve for x

OpenStudy (anonymous):

okay thank you much!

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

I hate to bother you, but could you help me with the rest that i have questions on please?